OpenStudy (anonymous):

find all solutions to 2sin2x-radical3=0

OpenStudy (adamaero):

well flip flop that equation around a bit!

OpenStudy (anonymous):

as in?

OpenStudy (adamaero):

try to think of it as a puzzle of pieces you already know: \[\sqrt{3}\] needs what to match something of a 30-60-90 triangle?

OpenStudy (anonymous):

so i would use the 60 degree triangle?

OpenStudy (adamaero):

see if the picture helps: http://www.themathpage.com/atrig/trig_IMG/270.gif

OpenStudy (anonymous):

alrighty! thank you! i appreciate it!

OpenStudy (anonymous):

could you help me zepdrix?

OpenStudy (adamaero):

Do you still need help on this one?

OpenStudy (anonymous):

haha yeah…

OpenStudy (adamaero):

Did you move the sqrt{3} to the other side, and divide by...

OpenStudy (anonymous):

then divide by 2

OpenStudy (anonymous):

yeah i did that

OpenStudy (anonymous):

then what identiy do i use for sin2x?

OpenStudy (adamaero):

\[\sin2x = \sqrt{3}/2\] and the part you might be caught on is the identity...right!

OpenStudy (anonymous):

yeah…thats the struggle.

OpenStudy (adamaero):

2sinxcosx = sin2x

OpenStudy (adamaero):

But it'd be easier just to know a property of the sinx function: http://upload.wikimedia.org/math/5/0/1/5014f43480bf84023f32b2d6059b0052.png

OpenStudy (adamaero):

pretend it's just sinx = \[\sqrt{3}/2\]

OpenStudy (anonymous):

right i got that

OpenStudy (adamaero):

so for that x would equal? Anyway, I have to go. ask people by putting @ before their name

OpenStudy (anonymous):

so would it just be pie/3+2pie n?

OpenStudy (anonymous):

ouh alrighty thanks!

OpenStudy (adamaero):

www.youtube.com/watch?v=z02AsqFuUnE